OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 0}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)

OpenStudy (jdoe0001):

hmm I should have said \(\bf y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\) but, there, find the slope and plug in the values

OpenStudy (skullpatrol):